OpenStudy (anonymous):

Let G be a group. Let H is a subset of  G. Then H is a subgroup of G if the following conditions are true (1) e belongs to H (where e is the identity of G). (2) if h, k belongs to H then hk belongs to H (H is closed). (3) if h belongs to H then h inverse belongs to H. what does it mean by h inverse???

OpenStudy (zarkon):

if \(h\in H\) then the inverse of \(h\) is an element \(h^{-1}\) such that \(h\times h^{-1}=e\)

OpenStudy (anonymous):

ok!!! thank you

OpenStudy (zzr0ck3r):

some examples of common inverses addition a+(-a) = 0 here 0 is the identity because x+0 = x so the additive inverse of a would be -a multiplication a*1/a = 1 here 1 is the identity because 1*x = x so the multiplicative inverse of a is a/a function composition f(x), then f inverse is the unique function such that f(g(x)) = x and g(f(x)) = x here f(x) = x is the identity because f(x) composed with x is just f(x)

OpenStudy (zzr0ck3r):

that should say "so the multiplicative inverse of a is 1/a"